OpenStudy (anonymous):

Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals.

OpenStudy (anonymous):

f(x) = 80x + 65 We know that 65 is our y intercept. So (0,65) is one point. We also know that 80 is our slope. From our equation, we know that every time x increases by one unit, y will increase by 80 units. So we also have (4,385) for example.

OpenStudy (anonymous):

All you have to do to graph something is find a minimum of two points, then draw a line between them.

OpenStudy (anonymous):

Do you get it Rocker?

OpenStudy (anonymous):

I'm piecing it together..

OpenStudy (anonymous):

f(x) = 80x + 65 Let x = 4 f(4) = 80(4) + 65 f(4) = 320 + 65 f(4) = 385 y = 385 x = 4 (4,385) You can let x be any value you want to find the corresponding value of y.

OpenStudy (anonymous):

So find two points and graph them?

OpenStudy (anonymous):

Yeah. :D

OpenStudy (anonymous):

And label the points inbetween?

OpenStudy (anonymous):

You only need two ordered pairs, because a line is infinite. If you labeled all the points on a line, you'd have to be stuck in limbo for all of eternity.

OpenStudy (anonymous):

So, how about x =4, and x= 6? xD then I get those ordered pairs and I put the pairs in as the answer?

OpenStudy (anonymous):

f(x) = 80x + 65 x = 4 Try solving it.

OpenStudy (anonymous):

f(x) = 80x + 65 f(4) = 80(4) + 65 f(4) = 320 + 65 f(4) = 385

OpenStudy (anonymous):

So, (4, 385)

OpenStudy (anonymous):

perfect :D

OpenStudy (anonymous):

now let x = 6 f(x) = 80x + 65

OpenStudy (anonymous):

f(x) = 80x + 65 f(6) = 80(6) + 65 f(6) = 480 + 65 f(6) = 545 (6, 545)

OpenStudy (anonymous):

You have mastered the f(x).

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

I understand how to find f(x) xD but the question has me in a whirl.

OpenStudy (anonymous):

Well just say that you have to solve for two points, and then show your work solving for x = 8 and x = 0. Just so you know, when you set x to zero, you'll find your y intercept.

OpenStudy (anonymous):

Okay. So.. My answer is the work for, and the two points, right?

OpenStudy (anonymous):

Yes, but make sure to describe each step as you do it, like 1. Solve for f(1) Which is just saying let x = 1.

OpenStudy (anonymous):

I'm assuming that 'interval' is the slope. Which you know already.

OpenStudy (anonymous):

Yes. lol

OpenStudy (anonymous):

Cool, four more right?

OpenStudy (anonymous):

So this is my answer for the whole section, right?

OpenStudy (anonymous):

And five more.

OpenStudy (anonymous):

You have to describe your steps in part C.

OpenStudy (anonymous):

Meaning?

OpenStudy (anonymous):

Explain why I'm doing what, as I'm doing it?

OpenStudy (anonymous):

Also, you don't let f(x) = 6, you solve for f(6).

OpenStudy (anonymous):

f(x) = 6 is it's own equation.

OpenStudy (anonymous):

Ohh, okay.

OpenStudy (anonymous):

Show me your result when it's done.

OpenStudy (anonymous):

What do you mean by describe

OpenStudy (anonymous):

"solve for f(0)"?

OpenStudy (anonymous):

Solve for f(x) = 80x + 65 when x = 4. f(4) = 80(4) + 65 [Substitute 4 for x] f(4) = 320 + 65 [Simplify] f(4) = 385 [Simplify] y = 385 [Substitute f(4) for y] (4, 385) [Graph point]

OpenStudy (anonymous):

Ahhh, okay. Haha,

OpenStudy (anonymous):

Yeah. :D

OpenStudy (anonymous):

Yay! Thanks, ready for the next?

OpenStudy (anonymous):

Post it. :)